A certain delivery service offers both express and standard delivery. Eighty percent of parcels are sent by standard delivery, and 20% are sent by express. Of those sent standard, 75% arrive the next day, and of those sent express, 94% arrive the next day. A record of a parcel delivery is chosen at random from the company's files.

a. What is the probability that the parcel was shipped express and arrived the next day?

Answer by mathmate(429)


Question:
A certain delivery service offers both express and standard delivery. Eighty percent of parcels are sent by standard delivery, and 20% are sent by express. Of those sent standard, 75% arrive the next day, and of those sent express, 94% arrive the next day. A record of a parcel delivery is chosen at random from the company's files.
a. What is the probability that the parcel was shipped express and arrived the next day?

Summary of given information:
standard delivery=80%
express=20%
Next day delivery for standard=75%
Next day deliver for express = 94%

Solution:
E=Event of shipping express
N=event of arriving next day
We look for the probability of a randomly selected parcel shipped express and arrived the next day. This is obtained by the multiplication of the two ratios, similar to a two-step experiment.
P(E∩N)=P(E)*P(N)=0.20*0.94=0.188